# main_arXiv.bib

### v1

@@ B.1 Proof of Lemma 1
$=\frac{\gamma}{\beta}\log\bigg{(}\frac{\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta X(s^{\prime})}}{\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta V(s^{\prime})}}\bigg{)}$
$=\frac{\gamma}{\beta}\log\bigg{(}\frac{\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta V(s^{\prime})+\beta(V(s^{\prime})-X(s^{\prime}))}}{\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta V(s^{\prime})}}\bigg{)}$
$\leq\frac{\gamma}{\beta}\log\bigg{(}\frac{\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta V(s^{\prime})+\beta\|V-X\|_{\infty}}}{\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta V(s^{\prime})}}\bigg{)}$
$=\frac{\gamma}{\beta}\log\bigg{(}e^{\beta\|V-X\|}\bigg{)}$
@@ B.3 Proof of Lemma 3
We report the proof in the case $\beta>0$, as the other case is proven using completely similar lines.
Since the policy $\pi$ is fixed, we omit it from the notation. Let $(s,a)$ be a state-action pair that satisfies $\|Q_{1}-Q_{2}\|=|Q_{1}(s,a)-Q_{2}(s,a)|$ and without loss of generality that $Q_{1}(s,a)\geq Q_{2}(s,a)$. Then,
@@
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\frac{\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}}{e^{-\beta\frac{1}{1-\gamma}}}$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}e^{\beta\frac{1}{1-\gamma}}\sum_{s^{\prime}}|P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)|$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}e^{\beta\frac{1}{1-\gamma}}\tau\,.$
@@
$\|Q_{1}-Q_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{\beta\frac{1}{1-\gamma}}}{\beta}\tau\,.$
Appendix C Lower Bound on Bernoulli Likelihood Ratio
In this section, we revisit and develop a technical result that bounds the likelihood ratio of two samples under different hypotheses on a high probability event. Parts of the proof closely resembles parts of Lemma 17 in [gheshlaghi2013minimax]; however, we stress that our treatment fixes an error in the proof, which however requires slightly stronger assumptions than those imposed in [gheshlaghi2013minimax]. In addition, while the result in [gheshlaghi2013minimax] only considers $p\geq\frac{1}{2}$, ours deal with both cases of $p\geq\frac{1}{2}$ and $p<\frac{1}{2}$.
@@ Appendix C Lower Bound on Bernoulli Likelihood Ratio
=$\mathcal{E}=\bigg{\{}\tilde{p}t-\tilde{k}\leq\sqrt{2p(1-p)\log(\frac{c_{2}}{2\theta})}\bigg{\}}\,,$
@@
For $c_{1}=32$, it holds that $\frac{L_{1}(W)}{L_{0}(W)}1_{\mathcal{E}}\geq\frac{2\theta}{c_{2}}1_{\mathcal{E}}$
@@
Define $m=t-k$, which is now the number of failed coin flips. Hence,
@@
Using this along with the fact that $(1-p)>p$ and $m\leq t$, we have
@@
Appendix D Proofs of Lower Bounds
D.1 Lower Bound for Q-value Learning
For a lower bound we let us inspire by the class of MDPs used in [azar2011reinforcement] where each MDP is of size $SA=3n$ for some $n\in\mathbb{N}$ and where each MDP consists of $n$ triplets of state-action pairs $\{z_{i}^{0},z_{i}^{G},z_{i}^{B}\}$ where the transition from $z_{i}^{0}$ to $z_{i}^{G}$ is given by $P(z_{i}^{G}|z_{i}^{0})=q_{i}$ and the transition from $z_{i}^{0}$ to $z_{i}^{B}$ is given by $P(z_{i}^{B}|z_{i}^{0})=1-q_{i}$. Both $z_{i}^{G}$ and $z_{i}^{B}$ are absorbing and the reward is $1$ in $z_{i}^{G}$ and 00 in both $z_{i}^{0}$ and $z_{i}^{B}$. For all three types of state-actions we can explicitly calculate the state-action value-functions
$Q(z_{i}^{0})$ $=\frac{-\gamma}{\beta}\log(q_{i}e^{-\beta\frac{1}{1-\gamma}}+1-q_{i})$
$Q(z_{i}^{G})$ $=\frac{1}{1-\gamma}$
$Q(z_{i}^{B})$ $=0$
for all $i=1,\ldots,n$. Denote the collection of all such MDPs by $\mathbb{M}$.
Fix a triplet with index $i$ and consider the two hypotheses $H^{i}_{0}:q_{i}=p$ and $H^{i}_{1}:q_{i}=p+\alpha$ where $p$ and $\alpha$ are given by
$p=\begin{cases}1-e^{-\beta\frac{1}{1-\gamma}}\quad&\text{for }\beta>0\\ e^{-|\beta|\frac{1}{1-\gamma}}\quad&\text{for }\beta<0\end{cases}$
@@ D.1 Lower Bound for Q-value Learning
$\alpha=8\varepsilon\frac{|\beta|}{\gamma}\frac{1}{e^{|\beta|\frac{1}{1-\gamma}}-1}$
for any $\varepsilon$ in the range
$\varepsilon<\frac{1}{40}\frac{\gamma}{|\beta|}(1-e^{-|\beta|})\,.$
We use $M_{0}$ to denote an MDP where $H_{0}^{i}$ holds for our fixed triplet, and $M_{1}$ to denote an MDP where instead $H_{0}^{i}$ holds and $\mathbb{E}_{0}$ and $\mathbb{P}_{0}$ as the expectations operator and probability operator under $H_{1}^{i}$ and similarly $\mathbb{E}_{1}$ and $\mathbb{P}_{1}$ under $H_{0}^{i}$. Fix any $(\varepsilon,\delta)$-correct $Q$-algorithm $\mathcal{U}$. We start by showing that with these parameter we have that $Q_{M_{1}}^{*}(z_{i}^{0})-Q^{*}_{M_{0}}(z_{i}^{0})>2\varepsilon$, which we do by casing on the sign of $\beta:$
@@
$Q^{*}_{M_{1}}(z_{i}^{0})-Q^{*}_{M_{0}}(z_{i}^{0})$ $=\frac{\gamma}{|\beta|}\log\bigg{(}\frac{(p+\alpha)e^{|\beta|\frac{1}{1-\gamma}}+1-p-\alpha}{pe^{|\beta|\frac{1}{1-\gamma}}+1-p}\bigg{)}$
@@
The case for $\beta>0$ is similar, although in this case we have $p=1-e^{-\beta\frac{1}{1-\gamma}}$ and use the inequality $\log(1+x)\leq x$ for all $x>-1$ to get that
$Q^{*}_{M_{1}}(z_{i}^{0})-Q^{*}_{M_{0}}(z_{i}^{0})$ $=-\frac{\gamma}{\beta}\log\bigg{(}\frac{(p+\alpha)e^{-\beta\frac{1}{1-\gamma}}+1-p-\alpha}{pe^{-\beta\frac{1}{1-\gamma}}+1-p}\bigg{)}$
@@
In particular this means that the events $B_{0}:=\{|Q^{*}_{M_{0}}(z_{i}^{0})-Q_{t}^{\mathcal{U}}(z_{i}^{0})|\leq\varepsilon\}$ and $B_{1}:=\{|Q^{*}_{M_{1}}(z_{i}^{0})-Q_{t}^{\mathcal{U}}(z_{i}^{0})|\leq\varepsilon\}$ are disjoint events Let $t$ be the number of times the algorithm tries $z_{i}^{0}$. Since $\mathcal{U}$ is $(\varepsilon,\delta)$-correct it holds that $\mathbb{P}_{0}(B_{0})\geq 1-\delta\geq\frac{3}{4}$.
Let $k$ be the number of transitions from $z_{i}^{0}$ to $z_{i}^{G}$ in the t trials. We then define $\tilde{k},\tilde{p}$ and $\theta$ by
@@
$\mathcal{E}$ =$=\bigg{\{}\tilde{p}t-\tilde{k}\leq\sqrt{2p(1-p)t\log(\frac{8}{2\theta})}\bigg{\}}$
for which we have $\mathbb{P}_{0}(\mathcal{E})>\frac{3}{4}$ by Lemma 16 in [gheshlaghi2013minimax] and thus $\mathbb{P}_{0}(B_{0}\cap\mathcal{E})>\frac{1}{2}$. Now by theorem 5 we get that
$\mathbb{P}_{1}(B_{0})\geq\mathbb{P}_{1}(B_{0}\cap\mathcal{E})=\mathbb{E}_{1}[1_{\mathcal{E}}1_{B_{0}}]=\mathbb{E}_{0}\bigg{[}\frac{L_{1}}{L_{0}}1_{\mathcal{E}}1_{B_{0}}\bigg{]}\geq\frac{\theta}{4}\mathbb{E}_{0}[1_{\mathcal{E}}1_{B_{0}}]=\frac{\theta}{4}\mathbb{P}_{0}(\mathcal{E}\cap B_{0})\geq\frac{\theta}{8}$
@@
$t<\frac{p(1-p)}{32\alpha^{2}}\log(\frac{1}{8\delta})$
@@
$\geq\frac{\gamma^{2}}{64\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}$
we conclude that if the algorithm $\mathcal{U}$ tries the state-action pair $z_{i}^{0}$ less than
@@
times under the hypothesis $H_{0}^{i}$ then $\mathbb{P}_{1}(B_{0})>\delta$ and $B_{0}\subset B_{1}^{c}$.
Next we use the fact that the structure of the MDPs is such that information on the $Q$-value of any state-action pair carries no information on the $Q$-values of any other state-action pair.
If the number of total transition samples is less than $\frac{n}{2}\tilde{T}(\varepsilon,\delta)=\frac{S}{6}\tilde{T}(\varepsilon,\delta)$ there must be at least one $n/2$ state-action pairs in the set $\{z_{i}^{0}\}_{i=1}^{n}$ that has been tried no more than $\tilde{T}(\varepsilon,\delta)$ times which without loss of generality we might assume are the state-action pairs $\{z_{i}^{0}\}_{i=1}^{n/2}$.
Let $T_{i}$ be the number of times the algorithm has tried $z_{i}^{0}$ for $i\leq n/2$ Due to the structure of the MDPs in $\mathbb{M}$ it is sufficient to consider only the algorithms that yields an estimate of $Q^{\mathcal{U}}_{T_{i}}$ based on samples from $z_{i}^{0}$ since any other samples can yield no information on $Q^{*}(z_{i}^{0})$
Thus by defining the events $\Lambda_{i}:=\{|Q_{M_{1}}^{*}(z_{i}^{0})-Q^{\mathcal{U}}_{T_{i}}(z_{i}^{0})|>\varepsilon\}$ we have that $\Lambda_{i}$ and $\Lambda_{j}$ are conditionally independent given $T_{i}$ and $T_{j}$. We then have
@@
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{n}/2=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}_{1}(\{T_{i}=t_{i}\}_{1\leq i\leq n/2})(1-\delta)^{n/2}$
@@
$\mathbb{P}_{1}(\{\Lambda_{i}^{c}\}_{1\leq i\leq n/2}|\{T_{i}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq n/2})\leq(1-\delta)^{\frac{n}{2}}$

### v2

@@ F.1 Proof of Lemma 1
$=-\frac{\gamma}{\beta}\log\bigg(\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta X(s^{\prime})-\beta(V(s^{\prime})-X(s^{\prime}))}\bigg)+\frac{\gamma}{\beta}\log\bigg(\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta X(s^{\prime})}\bigg)$
$\leq-\frac{\gamma}{\beta}\log\bigg(\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta X(s^{\prime})-\beta\|V-X\|}\bigg)+\frac{\gamma}{\beta}\log\bigg(\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta X(s^{\prime})}\bigg)$
@@ F.3 Proof of Lemma 3
There are four different cases to consider, namely the combinations arrising from $\beta>0$ vs $\beta<0$ and wether on the state-action pair $(s,a)$ that realizes the maximum it holds that $Q_{1}(s,a)>Q_{2}(s,a)$ or $Q_{2}(s,a)>Q_{1}(s,a)$.
Case 1: $\beta>0,Q_{1}(s,a)>Q_{2}(s,a)$.
@@
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\frac{|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}|}{e^{-\beta\frac{1}{1-\gamma}}}\,.$
@@
$\|Q_{1}-Q_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{\beta\frac{1}{1-\gamma}}}{\beta}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}\big|\,.$
Case 2: $\beta>0$ and $Q_{1}(s,a)<Q_{2}(s,a)$. The proof is very similar but the extension $V_{2}(s)=V_{1}(s)+V_{2}(s)-V_{1}(s)$ is now done in the numerator instead:
$\|Q_{1}-Q_{2}\|$ $=\frac{\gamma}{\beta}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{2}(s^{\prime})}}\bigg)$
$=\frac{\gamma}{\beta}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})-\beta(V_{2}(s^{\prime})-V_{1}(s^{\prime}))}}\bigg)$
$\leq\frac{\gamma}{\beta}\ln\bigg(e^{\beta\|V_{1}-V_{2}\|}\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$=\gamma\|V_{1}-V_{2}\|+\frac{\gamma}{\beta}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}-\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\frac{|\sum_{s^{\prime}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}|}{e^{-\beta\frac{1}{1-\gamma}}},$
which again yields
$\|Q_{1}-Q_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{\beta\frac{1}{1-\gamma}}}{\beta}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}\big|\,.$
Case 3: $\beta<0$ and $Q_{1}(s,a)>Q_{2}(s,a)$.
$\|Q_{1}-Q_{2}\|$ $=\frac{\gamma}{|\beta|}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{2}(s^{\prime})}}\bigg)$
$=\frac{\gamma}{|\beta|}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})-|\beta|(V_{2}(s^{\prime})-V_{1}(s^{\prime}))}}\bigg)$
$\leq\frac{\gamma}{|\beta|}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})-|\beta|\|V_{1}-V_{2}\|}}\bigg)$
$=\gamma\|V_{1}-V_{2}\|+\frac{\gamma}{|\beta|}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\ln\bigg(1+\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}-\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{|\beta|}\big|\sum_{s^{\prime}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{|\beta|V_{1}(s^{\prime})}\big|$
$\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{|\beta|}e^{|\beta|\frac{1}{1-\gamma}}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{|\beta|\big[V_{1}(s^{\prime})-\frac{1}{1-\gamma}\big]}\big|,$
which implies
$\|Q_{1}-Q_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{\beta}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{|\beta|[V_{1}(s^{\prime})-\frac{1}{1-\gamma}]}\big|.$
Case 4: $\beta<0$ and $Q_{2}(s,a)>Q_{1}(s,a)$. The proof of this case is similar to the other three cases and is omitted and the final part of the lemma follows by the triangle inequality and the fact that $e^{-|\beta|V_{1}(s)}\leq 1$ and $e^{-|\beta|(\frac{1}{1-\gamma}-V_{1}(s))}\leq 1$. ∎
Appendix G Lower Bound on Bernoulli Likelihood Ratio
In this section, we revisit and develop a technical result that bounds the likelihood ratio of two samples under different hypotheses on a high probability event. Parts of the proof closely resembles parts of Lemma 17 in [23]; however, we stress that our treatment fixes an error in the proof, which however requires slightly stronger assumptions than those imposed in [23]. In addition, while the result in [23] only considers $p\geq\frac{1}{2}$, ours deal with both cases of $p\geq\frac{1}{2}$ and $p<\frac{1}{2}$.
@@ Appendix G Lower Bound on Bernoulli Likelihood Ratio
$\mathcal{E}=\bigg\{\tilde{p}t-\tilde{k}\leq\sqrt{2p(1-p)\log(\frac{c_{2}}{2\theta})}\bigg\}\,$,
@@
For $c_{1}=32$, it holds that $\frac{L_{1}(W)}{L_{0}(W)}1_{\mathcal{E}}\geq\frac{2\theta}{c_{2}}1_{\mathcal{E}}$ .
@@
Define m=tkm=t-k, which is now the number of failed coin flips. Hence,
@@
Using this along with the fact that (1p)>p(1-p)>p and $m\leq t$, we have
@@
Appendix H Proofs of Lower Bounds
H.1 Lower Bound for Q-value Learning
For a lower bound we construct the following class of MDPs with $S^{\prime}:=S+2$ states and $A$ actions where the first states are labelled $S_{1},,...,s_{S},s^{G},s^{B}$ and the actions are labelled $a_{1},...,a_{A}$. The states $s^{G}$ and $s^{B}$ are absorbing under any actions and $R(s^{G},a)=1$ for all $j$ and $R(s^{B},a)=0$ for all $a\in A$. For the states $s\in\{s_{1},...,s_{S}\}$, we have that $R(s,a)=0$ for all $a\in A$. We have $SA$ state-action pair combinations from $\{s_{1},...,s_{S}\}\times A=:Z$ on which we assume some ordering allowing us to write $z_{i},i\in[SA]$. Finally for all state-action pairs $z_{i}\in[SA]$ we have $P(s^{G}|z_{i})=q_{i}$ and $P(s^{B}|z_{i})=1-q_{i}$ for some $q_{i}\in[0,1]$. The structure of this class of MDPs allows us to get lower bounds on the samples needed to learn the $Q$-value of each state-action pair $z_{i}$ and then use the fact that samples used to learn the $Q$-values for different state-action pairs bring no information on eachother to get the final bound.
$z_{i}$$s^{G}$$s^{B}$$q_{i}$$1-q_{i}$R=1R=1R=0R=0
Figure 1: Dynamics and rewards of the hard-to-learn MDP class
For any state-action pair we can explicitly calculate the state-action value-functions
$Q(z_{i})$ $=\frac{-\gamma}{\beta}\log(q_{i}e^{-\beta\frac{1}{1-\gamma}}+1-q_{i}),$
$Q(s^{G},a)$ $=\frac{1}{1-\gamma},$
$Q(s^{B},a)$ $=0.$
Denote the collection of all such MDPs by $\mathbb{M}$.
Fix any index $i$ and consider the two hypotheses $H^{i}_{0}:q_{i}=p$ and $H^{i}_{1}:q_{i}=p+\alpha$ where $p$ and $\alpha$ are given by
$p=\begin{cases}1-e^{-\beta\frac{1}{1-\gamma}}\quad&\text{for }\beta>0,\\ e^{-|\beta|\frac{1}{1-\gamma}}\quad&\text{for }\beta<0,\end{cases}$
@@ H.1 Lower Bound for Q-value Learning
$\alpha=8\varepsilon\frac{|\beta|}{\gamma}\frac{1}{e^{|\beta|\frac{1}{1-\gamma}}-1},$
for any $\varepsilon$ in the range $\varepsilon<\frac{1}{40}\frac{\gamma}{|\beta|}(1-e^{-|\beta|})$ .
We use $M_{0}$ to denote an MDP where $H_{0}^{i}$ holds and $M_{1}$ to denote an MDP where instead $H_{0}^{i}$ holds and $\mathbb{E}_{0}$ and $\mathbb{P}_{0}$ as the expectations operator and probability operator under $H_{1}^{i}$ and similarly $\mathbb{E}_{1}$ and $\mathbb{P}_{1}$ under $H_{0}^{i}$. Fix any $(\varepsilon,\delta)$-correct $Q$-algorithm $\mathcal{U}$. We start by showing that with these parameter we have that $Q_{M_{1}}^{*}(z_{i})-Q^{*}_{M_{0}}(z_{i})>2\varepsilon$, which we do by casing on the sign of $\beta:$
@@
$Q^{*}_{M_{1}}(z_{i})-Q^{*}_{M_{0}}(z_{i})$ $=\frac{\gamma}{|\beta|}\log\bigg(\frac{(p+\alpha)e^{|\beta|\frac{1}{1-\gamma}}+1-p-\alpha}{pe^{|\beta|\frac{1}{1-\gamma}}+1-p}\bigg)$
@@
The case for $\beta>0$ is similar, although in this case we have $p=1-e^{-\beta\frac{1}{1-\gamma}}$ and use the inequality $\log(1+x)\leq x$ for all x>1x>-1 to get that
$Q^{*}_{M_{1}}(z_{i})-Q^{*}_{M_{0}}(z_{i})$ $=-\frac{\gamma}{\beta}\log\bigg(\frac{(p+\alpha)e^{-\beta\frac{1}{1-\gamma}}+1-p-\alpha}{pe^{-\beta\frac{1}{1-\gamma}}+1-p}\bigg)$
@@
In particular, this means that the events $B_{0}:=\{|Q^{*}_{M_{0}}(z_{i})-Q_{t}^{\mathcal{U}}(z_{i})|\leq\varepsilon\}$ and $B_{1}:=\{|Q^{*}_{M_{1}}(z_{i})-Q_{t}^{\mathcal{U}}(z_{i})|\leq\varepsilon\}$ are disjoint events Let $t$ be the number of times the algorithm tries $z_{i}$. Since $\mathcal{U}$ is $(\varepsilon,\delta)$-correct it holds that $\mathbb{P}_{0}(B_{0})\geq 1-\delta\geq\frac{3}{4}$.
Let $k$ be the number of transitions from $z_{i}$ to $z_{i}^{G}$ in the t trials. We then define $\tilde{k},\tilde{p}$ and $\theta$ by
@@
$\mathcal{E}$ $=\bigg\{\tilde{p}t-\tilde{k}\leq\sqrt{2p(1-p)t\log(\frac{8}{2\theta})}\bigg\}$
for which we have $\mathbb{P}_{0}(\mathcal{E})>\frac{3}{4}$ by Lemma 16 in [23] and thus $\mathbb{P}_{0}(B_{0}\cap\mathcal{E})>\frac{1}{2}$. Now by Theorem 9 we get that
$\mathbb{P}_{1}(B_{0})\geq\mathbb{P}_{1}(B_{0}\cap\mathcal{E})=\mathbb{E}_{1}[1_{\mathcal{E}}1_{B_{0}}]=\mathbb{E}_{0}\bigg[\frac{L_{1}}{L_{0}}1_{\mathcal{E}}1_{B_{0}}\bigg]\geq\frac{\theta}{4}\mathbb{E}_{0}[1_{\mathcal{E}}1_{B_{0}}]=\frac{\theta}{4}\mathbb{P}_{0}(\mathcal{E}\cap B_{0})\geq\frac{\theta}{8}\,.$
@@
$t<\frac{p(1-p)}{32\alpha^{2}}\log(\frac{1}{8\delta}),$
@@
$\geq\frac{\gamma^{2}}{64\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}},$
we conclude that if the algorithm $\mathcal{U}$ tries the state-action pair $z_{i}$ less than
@@
times under the hypothesis $H_{0}^{i}$, then $\mathbb{P}_{1}(B_{0})>\delta$ and $B_{0}\subset B_{1}^{c}$.
Next we use the fact that the structure of the MDPs is such that information on the $Q$-value of any state-action pair in $Z$ carries no information on the $Q$-values of any other state-action pair in $Z$.
Let $n:=SA$. If the number of total transition samples is less than $\frac{n}{2}\tilde{T}(\varepsilon,\delta)$ there must be at least $n/2$ state-action pairs $z_{i}$ that has been tried no more than $\tilde{T}(\varepsilon,\delta)$ times which without loss of generality we might assume are the state-action pairs $\{z_{i}\}_{i=1}^{n/2}$.
Let $T_{i}$ be the number of times the algorithm has tried $z_{i}$ for $i\leq n/2$ Due to the structure of the MDPs in $\mathbb{M}$ it is sufficient to consider only the algorithms that outputs an estimate of $Q^{\mathcal{U}}_{T_{i}}$ based on samples from $z_{i}$ since any other samples can yield no information on $Q^{*}(z_{i})$
Thus by defining the events $\Lambda_{i}:=\{|Q_{M_{1}}^{*}(z_{i})-Q^{\mathcal{U}}_{T_{i}}(z_{i})|>\varepsilon\}$ we have that $\Lambda_{i}$ and $\Lambda_{j}$ are conditionally independent given $T_{i}$ and $T_{j}$. We then have
@@
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{n/2}=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}_{1}(\{T_{i}=t_{i}\}_{1\leq i\leq n/2})(1-\delta)^{n/2},$
@@
$\mathbb{P}_{1}(\{\Lambda_{i}^{c}\}_{1\leq i\leq n/2}|\{T_{i}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq n/2})\leq(1-\delta)^{\frac{n}{2}}\,.$